OpenStudy (anonymous):

I don't get this question: (-3)+(-18) divided by (+2) divided by (-3)= I don't get this question: (-3)+(-18) divided by (+2) divided by (-3)= @Mathematics

OpenStudy (saifoo.khan):

Samosas are cool.

OpenStudy (anonymous):

haha lol ya pic is sorta messed up

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

so can u help me with the ques?

OpenStudy (saifoo.khan):

Noway, i come here for fun.

OpenStudy (anonymous):

samosa sach mein achha dikh raha hai milega kya?????????

OpenStudy (anonymous):

saifoo bhuk tumko bhi lagi hogi kyun???

OpenStudy (anonymous):

wtf. .

OpenStudy (anonymous):

ftw....

OpenStudy (saifoo.khan):

lol sheg ismay koi shaq?

OpenStudy (saifoo.khan):

wtf = welcome to facebook.

OpenStudy (anonymous):

achha pehle jawab doge to jyada samosa milega kya

OpenStudy (saifoo.khan):

ftw = facebook the website.

OpenStudy (saifoo.khan):

lol @sheq.

OpenStudy (anonymous):

hahahahha thanks dear i too meant the same

OpenStudy (anonymous):

arey yar samosa hi gayab ho gaya

OpenStudy (anonymous):

dont mind was jokin

OpenStudy (anonymous):

wat?????????

OpenStudy (saifoo.khan):

lol sheg.

OpenStudy (saifoo.khan):

miss perfect, u got the solution?

OpenStudy (anonymous):

no!!! ughhh i dont get it!

OpenStudy (saifoo.khan):

where are u stuck?

OpenStudy (anonymous):

well first of all!, how do i start the equation?? i usually get this stuff but this one is confusing...

OpenStudy (saifoo.khan):

what are your teachers doing back at school?

OpenStudy (saifoo.khan):

Or what are u doing back at school when your teachers are teaching?

OpenStudy (anonymous):

lol i dont understand how they explain it! so how do i start it??? :/

OpenStudy (saifoo.khan):

Look, always remember the universal rule. PEDMAS. Parenthesis. Exponents Divide Multiply, Add Subtract. Always solve your equation like in this series.

OpenStudy (anonymous):

i use bedmas: brackets, exponents, division, multiplication, addition, and subtraction

OpenStudy (saifoo.khan):

same thing.

OpenStudy (anonymous):

so the problem is: (-3)+(-18) divided by (+2) divided by (-3) so i do . .

OpenStudy (saifoo.khan):

Right. so first thing, DIVIDE. -3 + (-18) DIVIDE by +2 divide -3 \[-3 + \frac{-18}{2} \div -3\] \[-3 +(-9) \div -3\] \[-3 + \frac{-9}{-3}\] \[-3 + 3\] 0

OpenStudy (anonymous):

so divide -18 and positive 2?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

=-9 and so then divide -9 by -3?

OpenStudy (saifoo.khan):

Correct.

OpenStudy (anonymous):

=3, and so that is where i am confused

OpenStudy (saifoo.khan):

So, now -3 + 3 equals?

OpenStudy (anonymous):

0

OpenStudy (saifoo.khan):

And 0 is the answer.

OpenStudy (anonymous):

o, wow, and i thought it was so much more complex. .. .lol

OpenStudy (anonymous):

gud work saifoo!!

OpenStudy (anonymous):

...

OpenStudy (saifoo.khan):

lol, now hit good answer please. -----> -----> ---->>> ^^^^^^^

OpenStudy (anonymous):

haha np :)

OpenStudy (saifoo.khan):

Thank you! :) i hope u got what i explained.

OpenStudy (anonymous):

can u help me with one more plz?

OpenStudy (anonymous):

and yes i did

OpenStudy (saifoo.khan):

Alright, post on left.

OpenStudy (anonymous):

arey saifoo dada kitne gold medal lene hai??

OpenStudy (anonymous):

k the problem is: 3x[14+(-18)] / (-4)

OpenStudy (anonymous):

use bodmas as explained by saifoo bhai

OpenStudy (anonymous):

...

OpenStudy (saifoo.khan):

use bedmas.

OpenStudy (anonymous):

ok.... so for the first step then the answer is -4?

OpenStudy (anonymous):

han wahi hai bodmas or bedmas

OpenStudy (saifoo.khan):

lol @sheq.

OpenStudy (anonymous):

....

OpenStudy (saifoo.khan):

YEs.

OpenStudy (anonymous):

ok then what? im cnfused...

OpenStudy (anonymous):

in place of over u writing exponential

OpenStudy (anonymous):

wat?

OpenStudy (saifoo.khan):

3x[-4] / (-4) \[\frac{3 \times -4}{-4}\] now look, -4 and -4 cancels out \[\frac{3 \times \cancel{-4}}{\cancel{-4}}\] Finally you are left with 3.

OpenStudy (anonymous):

ohh o kthankyou!

OpenStudy (anonymous):

so for this one : (-16) / 2x(3+1)

OpenStudy (anonymous):

it would be the 3+1 first right?

OpenStudy (saifoo.khan):

hmm.

OpenStudy (anonymous):

...

OpenStudy (saifoo.khan):

hmm = yes.

OpenStudy (anonymous):

oh ok lol then what?

OpenStudy (anonymous):

-16 / 2?

OpenStudy (anonymous):

wait i got the answer nvm! :D

OpenStudy (saifoo.khan):

Good to hear finally.

OpenStudy (anonymous):

lol, im not stupid, its just im nervous cuz im studying for a test . . tomorrow!

OpenStudy (saifoo.khan):

lol, hope u dont sleep in class.

OpenStudy (anonymous):

of course not! lol

OpenStudy (saifoo.khan):

lol, what grade r u in? that was so~ basic.

OpenStudy (anonymous):

well for u it was -_- 8 . . .. how old r u?

OpenStudy (saifoo.khan):

16. u?

OpenStudy (anonymous):

13

OpenStudy (saifoo.khan):

oh, cool.

OpenStudy (anonymous):

i dont get this one: 25+(-10)-3x[2-(-2)]=

OpenStudy (anonymous):

i have the +4 already

OpenStudy (saifoo.khan):

-4.

OpenStudy (anonymous):

what?

OpenStudy (saifoo.khan):

25+(-10)-3x[2-(-2)] 25+(-10)-3x[-4]=

OpenStudy (anonymous):

nvm got the answer! :)